OpenStudy (anonymous):

Sue's foot exerts a force of 66 N to the North and Jenny's foot exerts a force of 88 N to the East. What is the magnitude and direction of the resultant force on the ball?

OpenStudy (agreene):

 \[66^2+88^2 = c^2\] \[12100^{1/2}=c\] \[c=110N\] 45 degres NE